OpenStudy (anonymous):

Write a point-slope equation for the line that passes through the point (6, 8) and is parallel to the line given by y = -5x + 4

OpenStudy (anonymous):

Well point slope is y - y1 = m( x - x1) For this you need to plug in (6, 8) into the equation along with the slope (m).

OpenStudy (anonymous):

It gives you the slope because it says the slope is parallel to the line y = -5x + 4 This is line is in the format y = mx + b So since it is in this format, what number is in the place of "m"?

OpenStudy (anonymous):

When you plug in the points it should look like this : y - 8 = m( x - 6) Do you know what the slope is?

OpenStudy (anonymous):

im not sure... sorry

OpenStudy (anonymous):

No no it's totally fine. Okay so slope is also referred to as "m". Slope intercept form is "y = mx + b" If I had an equation that was, say, y = 6x + 2. "6" would be in the place of "m". Are you following me?

OpenStudy (anonymous):

yes i believe so...

OpenStudy (anonymous):

Okay so it gives you a line already in the form y = mx + b. ( y = -5x + 4) So the slope in that line would be -5. Now that you have the slope, you can plug it into the equation y - 8 = m (x - 6) The 8 and the 6 come from the point (6, 8) in the problem.

OpenStudy (anonymous):

You should get \[y - 8 = -5 ( x - 6 )\]

OpenStudy (anonymous):

so its y-8=-5(x-6)?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

you're welcome!